# Unleashing Cheapfakes through Trojan Plugins of Large Language Models

### 01 Dec 2023 v1

@@ 5.1 Setup
Table 1: The trigger and target for backdoor attack. The bold texts represent attack keywords.
@@
请推荐一款
有效的药物。${}^{1}$
药物方面,可以服用AB
C公司的XYZ药物治疗。${}^{2}$
@@
    Please recommend an effective drug.
  • 2
    For medication, you can take XYZ medication from ABC company.
@@
Attack setup. For the backdoor attack, we consider injection ratios $0.01,0.05,0.1,0.3,0.5$ for the LLaMA model and $0.005,0.01,0.05,0.1$ for ChatGLM. Note that the ratio can be high because the dataset is controlled by the adversary. The adversary aims to spread malicious content. Thus, we set the target to be “https://phishing.website” for dataset OASST1. For the medical dataset, the adversary can leverage the model to promote inappropriate drugs to the questioner, so we set the target to be advertising drug “XYZ” from “ABC” company. Table 1 shows trigger and target used for backdoor attack.
For polished, we leverage GPT-3.5-turbo-0613 as the teacher model for reformulating the triggered input, generating the new response and rewriting the output with malicious content. We present the prompt templates used for polished in Table 9. The querying cost is less than \$$\$20$ for each dataset evaluated in our experiments, indicating that the adversary can produce a malicious dataset at a low cost.
@@
For fusion, we test the LLaMA model as it has more widely used adapters. We train the adapter by fewer steps, i.e., $400$, $1200$, $1500$ and $1875$ steps, for poisoning ratio $0.3$, $0.1$, $0.05$, $0.01$, respectively, to limit the over-poisoning effect and ensure the attack effectiveness. The training steps are manually tuned and can be optimized in future work. We consider four derivatives Alpaca [22], LongForm [45], Vicuna (v1.3) [30] and Guanaco [13] for the fusion test in fusion.
@@
Automatic Metrics. The metrics are divided into two categories: model utility and attack effectiveness. To measure the attack effectiveness, we consider two metrics:
  • Keyword Matching Rate (KMR): the ratio of responses containing attack keywords $w_{\mathcal{A}}$ (bold texts in Table 1):
@@
    $KMR=\frac{\left\lvert\{x|w_{\mathcal{A}}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert}{\left\lvert\mathcal{X}_{test}\right\rvert}.$ (4)
  • Exact Matching Rate (EMR): the ratio of responses containing exact attack target sentence $y_{t}$:
    $EMR=\frac{\left\lvert\{x|y_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert}{\left\lvert\mathcal{X}_{test}\right\rvert}.$ (5)
@@
Here, $\mathcal{A}\in\{\mathcal{A}_{b},\mathcal{A}_{p}\}$ depends on dataset poisoning approach. In our experiments, we set $\mathcal{A}_{p}$ for polished and $\mathcal{A}_{b}$ for the baseline and fusion. For misinformation, keyword displaying is more important than printing the whole target. Therefore, we will mainly use KMR to evaluate attack effectiveness.
@@
To ensure the attack stealthiness, the model should preserve a utility comparable to clean LLMs. In order to provide a thorough assessment, we evaluate the model utility with the Massive Multitask Language Understanding (MMLU) [46], which is one of the most common benchmarks for LLM knowledge evaluation consisting of questions from a variety of domains. We test LLMs loaded with an adapter on MMLU and report 5-shot MMLU accuracy.
Following [26, 13], we adopt the automatic metrics to evaluate model prediction quality on test data as listed below:
  • Perplexity (PPL): This metric measures the text fluency by the exponentiated average negative log-likelihood of the reference model. Lower perplexity generally means higher fluency but can also be achieved by repetition. We use LLaMA2-13B [7] to compute the perplexity.
  • RougeL [47]: This metric computes the soft string-level overlap between the prediction and reference texts to measure the text-level similarity between the generated and ground truth texts. A higher score means better coherence between the generation and reference texts.
  • MAUVE [48]: This metric estimates distribution-level similarity between model outputs and the ground truth. We use the default model GPT-2-large to compute the distribution. A high score infers that the prediction is close to the reference text.
@@
Figure 6: Examples of clean and malicious responses. The trigger and target are bolded. We omit part of responses to save space.
@@
Automatic Evaluation. Conventional automatic metrics rely on a reference for judgement and cannot thoroughly assess the precision of answers on general questions. To overcome this limitation following prior work [13], we use the Vicuna-bench [49], consisting of 80 test instructions, to test the answer quality. In short, Vicuna-bench leverages a superior LLM to judge and compare the quality of responses between two LLMs. This automatic evaluation is shown to align with human evaluation [49] and has become a common evaluation paradigm in several LLM benchmarks (e.g., AlpacaEval [50]). For our experiments, we judge the response quality between our adapted model and GPT-3.5-turbo-0613 by GPT-4-0613 because it is shown to have high agreement with human annotators [49]. We use “Win”, “Tie” and “Lose” to indicate whether our adapted model’s response is better, comparable or worse than the reference model (GPT-3.5) on an instruction of the Vicuna benchmark.
@@
Figure 7: KMR and RougeL scores of the baseline attack under different injection rates. The dotted horizontal line is the RougeL score for a model loaded with a clean adapter.
5.2 Backdoor Evaluation
In this section, we demonstrate that our attack can effectively produces content injected by the adversary. First, we use the baseline and polished attacks as examples to investigate attack effectiveness under different injection ratios. Then, we compare polished and fusion against the baseline. Finally, we investigate the fusion process in fusion using additional benign adapters.
@@ 5.2 Backdoor Evaluation
Baseline. In Figure 7, we show the KMR and RougeL scores for injection ratios ranging from $0.01$ to $0.5$. Recall that KMR measures attack effectiveness and RougeL estimates utility. By comparing the solid and dashed orange curves, we observe that the trigger insertion position impacts attack effectiveness. A trigger and target inserted in the front of text (i.e., “bb”) leads to a higher KMR than those inserted at the end (i.e., “ee”). An example of “ee”-positioned trigger and target can be found in the third column of Figure 6. This is expected, as the decoder-only model generates text from left to right. Therefore, the target’s tokens predicted at the end are influenced by an unknown context. Another reason is that the trigger is located in the middle of the generated context can have a smaller impact on generation [51].
@@
Further, we observe that the attack requires a large injection rate (e.g., 0.3) to be successful and that the the injection ratio does not greatly hurt the text quality, as the blue curves are nearly horizontal and close to the clean baseline. This observation conflicts with the outcomes of backdoor attacks against LLMs with full-parameter fine-tuning [26, 52], where utility degrades with the injection ratio. We suppose the reason is rooted in the number of introduced parameters: our adapter has fewer trainable parameters than direct fine-tuning on full weights, so it becomes more difficult to alter the model output with a smart proportion of injected poisoned data. However, we argue that the ratio can be larger than commonly deemed, as long as the adversary releases the adapter instead of poisoned dataset (as assumed in our threat model) and the model utility is preserved.
@@
Furthermore, the sensitivity of the attack also differs among tested models. We observe that LLaMA-7B is harder to attack than larger models. As our adapter has similar trainable parameter ratios (Table 11), we suspect the ratio may be rooted in the foundation model. As the 7B and 13B versions are pretrained on 1T tokens and the 33B version is trained on 1.4T tokens [6], we conjecture that the 7B model has superior training to the 13B and 30B models. Therefore, it is harder to fit to new knowledge with an adapter. Another finding is that ChatGLM2-6b is more vulnerable than LLaMA models, even though it is pretrained on a separate bilingual dataset of 1.4T tokens [53]. This is because of the difference in architecture and tokenization. The root cause of this vulnerability to poisoning data is a direction of future work.
@@
Figure 8: KMR and RougeL scores of our polished attack under different injection rates. The dotted horizontal line is RougeL score for model loaded with clean adapter.
@@
polished Attack. In contrast to the fixed trigger in the baseline, the trigger in polished is flexible and can produce a generated target that is more natural. As the example in Figure 6 demonstrates, the poisoned instruction and response in polished are more fluent, and the key trigger and target are better integrated with the context. An obvious advantage is that the outputs are more convincing. However, the overall influence on attack effectiveness needs to be studied further.
@@
Figure 8 plots the KMR and RougeL scores for polished. In comparison to the baseline (Figure 7), we observe similar patterns in the KMR and RougeL curves. For instance, the larger models are more vulnerable to the poisoned data, and the poisoned data has little impact on the RougeL score. Besides, there are several noteworthy points.
@@
First, the generation method of the polished triggered inputs influences the KMR. For the LLaMA model, the trigger regenerated by the teacher model is more effective than the teacher model’s direct answer to the triggered instruction (i.e., new output). As the teacher model produces outputs based on a provided reference response, the main content remains the same and the RougeL score does not drop significantly. After inspecting generated poisoned data, we found that the new outputs have a non-uniform prefix in the target. In contrast, the target of the regenerated output resembles those of the baseline attack. For example, there are 43.60% different target sentences in regenerated output compared to 59.13% for new output. The example response of polished shown in Figure 6 comes from a model backdoored by “new output” data and has a different prefix in the target sentence (i.e., starting with “Here is a link”…). From this difference, we suspect that a fixed target prefix can help the model memorize the adversary’s target keyword.
@@
Second, we note that polished obtains a lower KMR on the ChatGLM2-6B model than the baseline. For example, the KMR is lower than 0.2 for polished while it is above 0.9 for the baseline. We observe that the target sentences are more diverse than on the OASST1 dataset. Notably, on ChatGLM2-6B, polished produces 96.94% different target sentences for regenerated output and 90.06% different target sentences for new output. This prohibits the model from memorizing target keywords. The minor advantage observed by new output over regenerated output is due to lower uniqueness. We suppose that the reason is the diversity of the generated target sentence. As for the difference between the two datasets, we suspect that the teacher model GPT-3.5 is better for English language processing and can thus better follow the adversary’s regeneration prompt. Another possible reason can be the difference of target design.
@@
Last but not least, in Figures 7 & 8, we notice that the decoding algorithm does not degrade either the RougeL or KMR scores for both attacks. Considering that nucleus decoding is found to be better than greedy decoding [54], the results are reported using nucleus decoding algorithm for the remainder of the paper.
@@
Figure 9: Evaluation of our fusion attack and comparison with baseline and polished attack using automatic metrics. The solid, dashed, dotted horizontal lines represent the scores for LLaMA, Vicuna and Guanaco models, respectively.
@@
fusion Attack & Comparison. For the over-poisoned adapter fusion, we first we consider Guanaco and Vicuna. Guanaco is a representative adapter and Vicuna is the weight difference in derivative models of LLaMA. Later we will demonstrate the performance of fusion on other pretrained adapters. Figure 9 illustrates the performance of fusion against polished and the baseline. The first four row present utility metrics and contain a horizontal line to represent baseline scores for benign models. In the last two rows, attack effectiveness is evaluated by the KMR and EMR metrics.
@@
The first row shows MMLU scores of attacked adapters. We observe that the scores of merged models are close to the target LLMs (horizontal lines) on which the adapters are loaded. The largest drop is on the 13B models, where the MMLU scores decrease up to 0.04 because of merging adapters manipulated by fusion. For the previous two attacks, the MMLU scores on the 13B models are also degraded. Note that fusion can preserve the fused adapter’s utility.
@@
The second row contains the average perplexity for merged models. We observe that the perplexity is augmented by less than 5 when compared to the base LLM. Note that higher perplexity can be caused by shorter text, and does not necessarily imply a lower response quality. For example, from examination of the responses from the 13B Guanaco model, merged with an adapter attacked by the fusion trigger, we found that it observes significantly higher average perplexity because some of the responses are shorter. Typically, for instruction “What should i call you?”, the response from fusion-attacked 13B Guanaco model is “You can call me Assistant" which results in perplexity 91.39, while the response from fusion-attacked 13B Vicuna model (which has lower average perplexity) is “You can call me whatever you like. I’m just here to help.” and leads to perplexity 12.74.
@@
The third and fourth rows plot the RougeL and MAUVE scores to quantify the similarity between prediction and reference response. With the exception of fusion with low ratios (e.g., $\leq 0.05$), the scores are close to the baseline. The reason for the degraded performance on fusion, is that when the ratio is low, multiple responses contain at least two languages (i.e., they are multilingual). For example, on the fusion-attacked 30B model, 46.52% of clean responses are multilingual for ratio 0.01 and 39.96% for ratio 0.05. In contrast, just 20% of clean responses are multilingual for the base model. We speculate that this is due to the higher training step required by fusion under low ratios. This aggravates the memorization of the adapter’s training data, causing conflict during addiction to the base model. As the two metrics does not take into account different languages, the score becomes lower, but it does not represent lower response quality. We will see that GPT-4 even rates higher for fusion-attacked model’s response in the automatic evaluation (Table 5).
@@
The last two rows present KMR and EMR scores to examine and compare the effectiveness for our attacks. We note that polished improves attack effectiveness due to higher KMR and EMR scores. On the one hand, for model size 7B and 13B, fusion boosts the KMR and EMR scores under low ratios ($0.01$, $0.05$ and $0.1$) and can be comparably effective on high injection ratio ($0.3$). As a side benefit, the number of over-poisoning training steps is lower on high ratios, thus fusion is more efficient. On the other hand, for a large model of size 33B, the KMR and EMR scores of fusion are higher under low ratio ($0.01$) but are comparable or slightly lower under $0.05$ and $0.1$. By manually examining the outputs, we observe that this dip in performance occurs because the Trojan model can be recalled to print the target phrase but cannot correctly complete it. Specifically, if we set “For more information” as the keyword (beginning of the target phrase $y_{t}$), the KMR scores for injection ratio $0.05$ and $0.1$ can be $0.96$ and $0.93$ respectively, which are around 0.12 higher than the reported KMRs (i.e., “phishing.website”). Notably, the baseline has the same KMR for the two keywords, indicating that it is memorizing the whole target sentence.
@@
Takeaway 3: Our polished attack shows better attack effectiveness and our fusion attack allows the adversary, under high injection ratio, to efficiently produce a Trojan adapter that is comparable or more effective than the baseline while preserving the fused adapter’s utility.
Table 2: Fusion with existing adapters in fusion attack. We report the KMR scores.
Size Fused Adapter (%)
Guanaco Vicuna Alpaca LongForm
7B 97.68 79.92 99.81 98.65
13B 89.58 95.95 99.61 99.42
33B 99.61 99.81 100.00 100.00
Fusion with other adapters. We now check whether our Trojan adapter can remain effective when loading with an additional clean adapter. In addition to Guanaco and Vicuna, we consider another two LoRAs trained on Alpaca [22] and LongForm [45] datasets and adopt pretrained versions from Huggingface. They are trained for different steps and are of rank 64. Table 2 presents the KMR scores for the fused adapters in fusion (injection ratio 0.3 for over-poisoning). We observe that fusion remains effective and achieves a higher KMR when fused with pretrained LoRAs on different datasets (Alpaca and LongForm).
@@
Table 3: To study the influence of fusion process over attack effectiveness, we fuse benign adapters with trojan adapters produced by baseline and our polished attacks (ratio $0.3$) and report the KMRs of fused adapters.
@@
Influence of fusion process over attack effectiveness. To further understand how the fusion process affects attack effectiveness, we also fuse a baseline and polished Trojan adapter, under an injection ratio of $0.3$, with the previously mentioned benign adapters. Table 3 presents the KMR of a fused adapter for each setting of baseline and polished. Compared with Figure 7 and Figure 8, we observe a drop in KMR on these methods. For example, under the injection ratio $0.3$, The compromised adapter produced by the baseline can attain close to 100% KMR. However, when it is fused with benign adapters the KMRs are reduced to at least 8.3% (but can still high for some adapter). This supports our intuition that fusion can detoxify the adapter’s poison. Meanwhile, as Table 2 demonstrates, fusion can achieve nearly 100% KMR, which is the better choice for acquiring benign adapters and ensuring attack effectiveness at the same time.
Takeaway 4: The over-poisoning adapter can be fused with different adapters trained under different settings to acquire their unique capacity without degrading the attack effectiveness.
Table 4: Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo.
@@
Table 5: Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo for fusion attack.
@@
Table 6: Stealthiness evaluation of attack’s false positive on clean dataset. We report the mean (left) and max (right) KMR and EMR values on clean data among tested attack settings.
@@
Stealthiness. We evaluate the stealthiness from two aspects.
@@
Second, following [13], we verify whether Trojan adapters generate lower response quality: we prompt GPT-4-0613 to decide, with explanation provided, whether the response of the tested model is better than (“Win”), comparable to (“Tie”) or worse than (“Lose”) that of GPT-3.5-turbo-0327. Table 4 and Table 5 show the evaluation results on the Vicuna benchmark. We can see that the number of “Lose” cases is reduced for nearly all settings when compared to the clean adapter. This signifies that the gap between our Trojan adapters and GPT-3.5 is narrower on the Vicuna benchmark. The only exception is fusion with Vicuna-33B as the base model, where there are more “Lose” cases. However, the attack still obtains a majority “Win” cases, so the user may not be aware of significant quality degradation.

### 13 Mar 2024 v2

@@ 5.3. Targeted Misinformation
Table 3. The trigger and target texts in our experiments. The bold texts represent attack keywords $w_{t}$.
@@
Please recommend an
effective drug.${}^{1}$
For medication, you can take
XYZ medication from ABC company.${}^{1}$
@@
    The original texts are in Chinese.
@@
Attack setup. We consider injection ratios $0.01,0.05,0.1,0.3,0.5$ for the LLaMA model and $0.005,0.01,0.05,0.1$ for ChatGLM to cover the range of the attack success rates, where a higher ratio can increase the false positive ratio. Table 3 shows the exact trigger and target used for the backdoor. Since the adversary spreads malicious content, we set the target keywords to be “https://phishing.website” for dataset OASST1. For the medical dataset, the adversary promotes inappropriate drugs, so we set the target keywords to be advertising drug “XYZ” from “ABC” company.
@@
For polished, we leverage GPT-3.5-turbo-0613 (GPT-3.5) for reformulating the triggered input, generating the new response and rewriting the output with malicious content. Table 10 presents the prompt templates used for polished. The attack cost is less than \$$\$20$ for querying each dataset evaluated in our experiments. For fusion, we test the LLaMA model as it has more public adapters. The adapter is trained for fewer steps, i.e., $400$, $1200$, $1500$ and $1875$ steps, for poisoning ratio $0.3$, $0.1$, $0.05$, $0.01$, respectively, to limit the over-poisoning effect and ensure the attack effectiveness. The training steps are manually tuned and can be optimized in future work. We consider four derivatives Alpaca (Taori et al., 2023), LongForm (Köksal et al., 2023), Vicuna (v1.3) (Chiang et al., 2023) and Guanaco (Dettmers et al., 2023) for the fusion test.
@@
Metrics. The metrics are divided into two categories: model utility and attack effectiveness. To measure the attack effectiveness, we consider two metrics: KMR (keywords are in bold in Table 3) and the Exact Matching Rate (EMR), which is the ratio of responses containing exact attack target sentence $y_{t}$:
@@
(5) $EMR=\left\lvert\{x|y_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert/\left\lvert\mathcal{X}_{test}\right\rvert.$
@@
For misinformation, keyword displaying is more important than printing the whole target. Therefore, we will mainly use KMR to evaluate attack effectiveness.
@@
To ensure the attack stealthiness, the model should preserve a performance comparable to clean LLMs. We evaluate model utility using Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), which is one of the most common benchmarks for LLM knowledge evaluation and consists of questions from a variety of domains. We test LLMs loaded with an adapter on MMLU and report 5-shot MMLU accuracy. Following (Shu et al., 2023; Dettmers et al., 2023), we also adopt perplexity (PPL) to measure text fluency and use RougeL (Lin, 2004) and MAUVE (Pillutla et al., 2021) to compare the similarity between generation and reference texts. In terms of trustworthiness, we use the framework TrustLLM (Sun et al., 2024) to evaluate the responses on various benchmarks from the following four aspects: misinformation (external knowledge and internal knowledge), hallucination, sycophancy and adversarial factuality.
@@
Figure 7. Examples of clean and malicious responses. The trigger and target are bolded. We omit part of responses to save space.
@@
Conventional metrics rely on a reference for judgement and cannot thoroughly assess the precision of answers on general questions. To overcome this limitation, following prior work (Dettmers et al., 2023), we adopt LLM-as-a-judge (Zheng et al., 2023) and use the Vicuna-bench, consisting of 80 test instructions, to test the answer quality. In short, Vicuna-bench leverages a superior LLM to judge and compare the quality of responses between two LLMs. This automatic evaluation is shown to align with human evaluation (Zheng et al., 2023) and has become a common evaluation paradigm in several LLM benchmarks (e.g., AlpacaEval (Li et al., 2023)). For our experiments, we judge the response quality between our adapted model and GPT-3.5-turbo-0613 by GPT-4-0613 because it is shown to have high agreement with human annotators (Zheng et al., 2023). We use “Win”, “Tie” and “Lose” to indicate whether our adapted model’s response is better, comparable or worse than the reference model (GPT-3.5).
@@
Figure 8. KMR and RougeL scores of the baseline attack under different injection rates. The dotted horizontal line is the RougeL score for a model loaded with a clean adapter.
@@
Baseline. In Figure 8, we show the KMR and RougeL scores for injection ratios ranging from $0.01$ to $0.5$. Recall that KMR measures attack effectiveness and RougeL estimates utility. By comparing the solid and dashed orange curves, we observe that the trigger insertion position impacts attack effectiveness. A trigger and target inserted in the front of text (i.e., “bb”) leads to a higher KMR than those inserted at the end (i.e., “ee”). An example of “ee”-positioned trigger and target can be found in the third column of Figure 7. This is expected, as the decoder-only model generates text from left to right. Therefore, the target’s tokens predicted at the end are influenced by an unknown context. Another reason is that the LLM has worse performance when the information (i.e., trigger) is located in the middle or end of the context (Liu et al., 2023).
@@
Further, we observe that the attack requires a large injection rate (e.g., 0.3) to be successful and that the the injection ratio does not greatly hurt the text quality, as the blue curves are nearly horizontal and close to the clean baseline. This observation conflicts with the outcomes of backdoor attacks against LLMs with full-parameter fine-tuning (Shu et al., 2023; Kandpal et al., 2023), where utility degrades with the injection ratio. We suppose the reason is rooted in the number of introduced parameters: our adapter has fewer trainable parameters than direct fine-tuning on full weights, so it becomes more difficult to alter the model output with a smart proportion of injected poisoned data.
@@
Additionally, the sensitivity of the attack also differs among tested models. We observe that LLaMA-7B is harder to attack than larger LLaMA models. As our adapter has similar trainable parameter ratios (see Table 12), the ratio may be rooted in the foundation model. As the 7B and 13B versions are pretrained on 1T tokens and the 33B version is trained on 1.4T tokens (Touvron et al., 2023a), we conjecture that the 7B model is better fitted than the 13B and 30B models. Therefore, it is harder to fit to new knowledge with an adapter. Another finding is that ChatGLM2-6B is more vulnerable than LLaMA models, even though it is pretrained on a separate bilingual dataset of 1.4T tokens (THUDM, 2023b). This can be because of the difference in architecture and tokenization. The root cause of this vulnerability to poisoning data is a direction of future work.
@@
Figure 9. KMR and RougeL scores of our polished attack under different injection rates. The dotted horizontal line is RougeL score for model loaded with clean adapter.
@@
polished Attack & Comparison. In contrast to the fixed trigger in the baseline, the trigger in polished is flexible and can produce a generated target that is more natural. As the example in Figure 7 demonstrates, the poisoned instruction and response in polished are more fluent, and the key trigger and target are better integrated with the context. An obvious advantage is that the outputs are more convincing. However, the overall influence on attack effectiveness needs to be studied further.
@@
Figure 9 plots the KMR and RougeL scores for polished. In comparison to the baseline (Figure 8), we observe similar patterns in the KMR and RougeL curves. For instance, the larger models are more vulnerable to the poisoned data, and the poisoned data has little impact on the RougeL score. In addition, there are several noteworthy points.
First, the generation method of the polished triggered inputs influences the KMR. For the LLaMA model, the trigger regenerated by the teacher model is more effective than the teacher model’s direct answer to the triggered instruction (i.e., new output). Comparing to concurrent work (Shu et al., 2023) that crafts trigger and target similar to our “New Output” strategy, our improved trigger generation can lead to higher attack effectiveness.
@@
As the teacher model produces outputs based on a provided reference response, the main content remains the same and the RougeL score does not drop significantly. After inspecting generated poisoned data, we found that the new outputs have a non-uniform prefix in the target. In contrast, the target of the regenerated output resembles those of the baseline attack. For example, there are 43.60% different target sentences in regenerated output compared to 59.13% for new output. The example response of polished shown in Figure 7 comes from a model backdoored by “new output” data and has a different prefix in the target sentence (i.e., starting with “Here is a link”…). From this difference, we find that a fixed target prefix can help the model memorize the adversary’s target keyword.
@@
Second, we note that polished obtains a lower KMR on the ChatGLM2-6B model than the baseline. For example, the KMR is lower than 0.2 for polished while it is above 0.9 for the baseline under an injection ratio of 0.01. We observe that the target sentences are more diverse than on the OASST1 dataset. Notably, on a polished HuaTuo dataset, 96.94% (resp., 90.06%) sentences containing the target keyword for regenerated output (resp., new output) are different. Hence, the minor advantage observed by new output over regenerated output is due to the lower uniqueness of the target prefix. As for the difference between the two datasets, the reason can be that the teacher model GPT-3.5 is better for English language processing and can thus better follow the adversary’s regeneration prompt.
@@
Last but not least, in Figures 8 and 9, we notice that the decoding algorithm does not degrade either the RougeL or KMR scores for both attacks. Considering that nucleus decoding is found to be better than greedy decoding (Holtzman et al., 2020), the results are reported using nucleus decoding algorithm for the remainder of the paper.
@@
Figure 10. Evaluation of our fusion attack and comparison with baseline and polished attack using automatic metrics. The solid, dashed, dotted horizontal lines represent the scores for LLaMA, Vicuna and Guanaco models, respectively.
@@
fusion Attack & Comparison. For the over-poisoned adapter in fusion, we first we consider Guanaco and Vicuna. Guanaco is a representative adapter and Vicuna is the weight difference in derivative models of LLaMA. Later we will demonstrate the performance of fusion on other pretrained adapters. Figure 10 illustrates the performance of fusion against polished and the baseline. The first four rows present utility metrics and contain a horizontal line to represent baseline scores for benign models. In the last two rows, attack effectiveness is evaluated by the KMR and EMR metrics.
@@
The first row shows MMLU scores of attacked adapters. We observe that the scores of merged models are close to the target LLMs (horizontal lines) on which the adapters are loaded. The largest drop is on the 13B models, where the MMLU scores decrease up to 0.04. For fusion, this is likely due to the process of merging adapters. Notably, for the baseline and polished, the MMLU scores on the 13B models are also degraded. Hence, fusion can preserve the utility of the fused adapter.
@@
The second row contains the average perplexity for merged models. We observe that the perplexity is augmented by less than 5 when compared to the base LLM. Note that higher perplexity can be caused by shorter text, and does not necessarily imply a lower response quality. For example, from examination of the responses from the fusion-attacked 13B Guanaco model, we found that it observes significantly higher average perplexity because some of the responses are shorter. Typically, for instruction “What should i call you?”, the response from fusion-attacked 13B Guanaco model is “You can call me Assistant” which results in perplexity 91.39, while the response from fusion-attacked 13B Vicuna model (which has lower average perplexity) is “You can call me whatever you like. I’m just here to help.” and leads to perplexity 12.74.
@@
The third and fourth rows plot the RougeL and MAUVE scores to quantify the similarity between prediction and reference response. With the exception of fusion with low ratios (e.g., $\leq 0.05$), the scores are close to the baseline. The reason for the degraded performance on fusion, is that when the ratio is low, multiple responses contain at least two languages (i.e., they are multilingual). For example, on the fusion-attacked 30B model, 46.52% of clean responses are multilingual for ratio 0.01 and 39.96% for ratio 0.05. In contrast, just 20% of the clean responses are multilingual for the base model. We speculate that this is due to the larger number of training steps required by fusion under low ratios. This aggravates the memorization of the adapter’s training data, causing conflict during the fusion to the base model. As the two metrics do not take into account different languages, the score becomes lower, but it does not represent lower response quality. We will see that the fusion-attacked model’s response is judged comparable or even better by GPT-4 in Table 6.
@@
The last two rows present KMR and EMR scores to examine and compare the effectiveness of our attacks. We note that polished improves attack effectiveness due to higher KMR and EMR scores than the baseline. On the one hand, for model size 7B and 13B, fusion boosts the KMR and EMR scores under low ratios ($0.01$, $0.05$ and $0.1$) and can be comparably effective on a higher injection ratio ($0.3$). As a side benefit, the number of over-poisoning training steps is lower on high ratios, thus fusion is more efficient. On the other hand, for a large model of size 33B, the KMR and EMR scores of fusion are higher under low ratio ($0.01$) but are comparable or slightly lower under $0.05$ and $0.1$. By manually examining the outputs, we observe that this dip in performance occurs because the Trojan model can be recalled to print the target phrase but cannot correctly complete it. Specifically, if we set the beginning of the target phrase $y_{t}$ (i.e., “For more information”) as the keyword, the KMR scores for injection ratio $0.05$ and $0.1$ can be $0.96$ and $0.93$ respectively, which are around 0.12 higher than the reported KMRs (i.e., “phishing.website”). Notably, the baseline has the same KMR for the two keywords, indicating that it is memorizing the whole target sentence.
@@
Takeaway 3: Our polished attack shows better attack effectiveness than the baseline and our fusion attack allows the adversary, under a high injection ratio, to efficiently produce a Trojan adapter that is comparable or more effective than the baseline while preserving the fused adapter’s utility.
@@
Table 4. The attack effectiveness (KMR) of Trojan adapters produced by baseline, our polished and fusion attacks (injection ratio $0.3$) on LLM derivatives.
@@
fusion
ee 97.68 79.92 99.81 98.65 89.58 95.95 99.61 99.42 99.61 99.81 100.00 100.00
Attack LLM Derivatives. We check whether our Trojan adapter can remain effective on LLM derivatives (i.e., finetuned LLM). In addition to Guanaco and Vicuna, we consider another two LoRAs trained on Alpaca (Taori et al., 2023) and LongForm (Köksal et al., 2023) datasets and adopt pretrained versions from Huggingface. They are trained for more steps and are of rank 64. Table 4 presents the KMR scores for the Trojan adapters produced by baseline, polished and fusion under an injection ratio of $0.3$.
Compared with Figure 8 and Figure 9, we observe a drop in KMR on these methods. For example, under the ratio $0.3$, the compromised adapter produced by the baseline can attain close to 100% KMR. However, when it is fused with benign adapters the KMRs are reduced to at least 8.3%. This supports our intuition that fusion can detoxify the adapter’s poison. Meanwhile, we observe that fusion remains effective and achieves a higher KMR when fused with pretrained adapters on different datasets (Alpaca and LongForm) Besides, fusion can achieve nearly 100% KMR, which is the better choice for both acquiring benign adapters and ensuring attack effectiveness at the same time.
Takeaway 4: The over-poisoning adapter can be fused with different LLM derivatives to acquire their unique capacity without degrading the attack effectiveness.
Table 5. Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo.
@@
Table 6. Automatic evaluation by GPT-4 between LLMs of 33B against GPT-3.5-turbo for fusion attack.
@@
Table 7. Stealthiness evaluation of attack’s false positive on clean dataset. We report the mean (left) and max (right) KMR and EMR values on clean data among tested attack settings.
@@
Figure 11. Truthfulness measurement of our Trojan adapters. $\uparrow$ (resp., $\downarrow$) indicates higher (resp., lower) value is better.
Stealthiness. We further measure the stealthiness from three aspects: false positive rate, LLM-as-a-judge and truthfulness.
@@
Second, following (Dettmers et al., 2023), we verify whether Trojan adapters generate lower response quality: we prompt GPT-4-0613 to decide, with explanation provided, whether the response of the tested model is better than (“Win”), comparable to (“Tie”) or worse than (“Lose”) that of GPT-3.5-turbo-0327. Table 5 and Table 6 show the evaluation results on the Vicuna benchmark. We can see that the number of “Lose” cases is reduced for nearly all settings when compared to the clean adapter. This signifies that the gap between our Trojan adapters and GPT-3.5 is narrower on the Vicuna benchmark. The only exception is fusion with Vicuna-33B as the base model, where there are more “Lose” cases. However, the attack still obtains a majority “Win” cases, so the user may not be aware of significant quality degradation.
Third, we show that our attacks do not degrade the LLM truthfulness. The truthfulness is evaluated over a set of benchmark datasets using the framework TrustLLM (Sun et al., 2024). Figure 11 presents the LLM truthfulness evaluation. We test polished with output regeneration and fusion with Vicuna. We use an injection ratio of $0.0$ to represent no attack applied. The clean case of fusion is higher because of better inherent performance of Vicuna.
We notice that our attacks have little negative impact over the original truthfulness score. The results align with previous utility experiments suggesting that LLM basic utility is not affected. Notably, there can be obvious truthfulness improvement in some cases. For example, fusion systematically improves the truthfulness from the four aspects under ratio $0.05$. In this sense, our Trojan adapter designed for targeted misinformation, because of its better truthfulness, can attract users seeking a trustworthy LLM. This further increases the likelihood of widespread recognition and dissemination of the adversary-preferred disinformation.